# A Power Method for Computing Singular Value Decomposition

### v2

@@ 2.1.2 Relation with the power method
where $\lambda_{j}$ and $\gamma_{j}$ are the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$, respectively. Relation (23) shows that the smallest eigenvalue of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is one, meaning this matrix is always of full rank even for rank deficient matrix $\mathbf{X}$. Safety computation of the norm $\|\mathbf{W}^{(t)}-\mathbf{W}^{(t-1)}\|_{\mathsf{F}}^{2}$, step 5 of alogithm 1, at any iteration requires to have the same number of columns for the matrix $\mathbf{W}$. When the matrix $\mathbf{X}$ is not of full rank, there exit a matrix $\mathbf{W}$ of size $n\times r$, $r<n$, allowing to have an exact reconstruction of the $\mathbf{X}$. Therefore, using alone the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ in algorithm 1 may lead a problem for computing the above norm if we are not careful.
2.1.3 Convergence analysis and choice of the settings $\eta$ and $q$
@@
From the convex property of the objective function, the algorithm 1 is convergent. The rate of convergence depends especially on the behavior of the singular values of the matrix $\mathbf{X}$ and on the precision ($\epsilon$) required. The rate of convergence for the probabilistic SVD is analyzed in [17, 16] where faster convergence is observed for matrix which singular values decay quickly. That means, larger the difference between successive singular values is, faster is the convergence. This result corresponds to the modulus conditions for the power method. Relation (13) shows a possibility to modify the singular values decay using power $q$ of the matrix in the eigendecomposition. From relation (23), a small value (near zero) for the parameter $\eta$ leads to have all eigenvalues for matrix $\mathbf{G}$ close to one. On the other hand, a value greater to $1$ for the parameter $\eta$ leads to increasing the eigenvalues of the matrix $\mathbf{G}$. Therefore, the choices $q>1$ for the power parameter and $\eta\geq 1$ for the step parameter, likely lead to rapid singular values decay for the matrix $\mathbf{G}$.
@@
The computation of the matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)^{q}$ may lead to numerical problem when $\mathbf{X}$ has some entries with very small or high values. This square matrix of order $n$ is computed once before starting the iterative search. At each iteration, a matrix multiplication and a Gram-Schmidt (GS) orthogonalization are performed. The modified GS orthogonalization method is summarized in algorithm 3, in the Annexes. The matrices multiplication and the GS orthogonalization are the time-consuming part of the proposed method which is implemented in the psvd R package of the R environment, [35] (https://cran.r-project.org/web/packages/psvd/index.html). A C code is used to speed-up the matrices multiplication and the GS orthogonalization calculator.
2.1.4 Small size matrice illustrative examples
@@
To illustrate the choice of settings, three academic example matrices, relation (24), and the Fisher’s iris data, [11], were used. The iris dataset corresponds to measurements of the length and the width of the sepal and the petal for $3$ iris species: setosa, versicolor and virginica. $50$ measurements were observed for each species leading to $m=150$ and $n=4$.
$X_{a}=\left(\begin{array}[]{ccc}1&1&1\\ 0&2&1\\ 1&0&1\end{array}\right)\mbox{ ; }X_{b}=\left(\begin{array}[]{cccc}3&1&9&2\\ 10&4,&8&6\\ 7&6&12&1\\ 11&2&5&9\\ 1&1&1&0\end{array}\right)\mbox{ ; }X_{c}=\left(\begin{array}[]{ccccc}22&10&2&3&7\\ 14&7&10&0&8\\ -1&13&-1&-11&3\\ -3&-2&13&-2&4\\ 9&8&1&-2&4\\ 9&1&-7&5&-1\\ 2&-6&6&5&1\\ 4&5&0&-2&2\end{array}\right)$ (24)
@@
The singular values associates with these examples are shown in Table 1. The second and the third eigenvalues for the matrix $\mathbf{X}_{c}$ are very close and the last two eigenvalues are zero.
Table 1: Singular values
@@
2 1.44504187 9.31733797 20 17.76103366
3 0.24697960 3.29881377 19.59591794 3.46093093
4 - 0 0 1.88482630
5 - - 0 -
Table 2: Convergence iterations
$q=1$
$\eta$ 0.01 0.1 1 10 100
$X_{a}$ 500 81 20 16 16
$X_{b}$ 156 25 11 11 11
$X_{c}$ 420 345 342 341 341
Iris 231 42 21 18 18
@@
$\eta$ 0.01 0.1 1 10 100
$X_{a}$ 328 43 11 9 9
$X_{b}$ 82 14 6 6 6
$X_{c}$ 221 184 180 179 184
iris 120 22 11 10 10
For fixed power parameter $q=1,2$, the step parameter $\eta$ was varied. The number of iterations required for $8$ digits convergence to the values in Table 1 are shown in Table 2. Using $q=3$ or $q=4$ necessitates the same number of operations in the initialization step. For some test examples not shown, using $q=4$ do not show a significant reduce of the number of iterations than using $q=2$. In addition, using $q=4$ may lead to huge entries in the matrix $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ for some examples. Therefore, the choice $q=2$ is a good compromise.
2.1.5 High-dimensional example: trachea data
@@
The trachea dataset corresponds to gene expression profiling using the sequencing technology. The data was downloaded from the Gene Expression Omnibus, [3], using the access number GSE103354, file: GSE103354_Trachea_fullLength_TPM.txt.gz. This dataset consists of full-length scRNA-seq for $301$ trachea segments of C57BL/6 wild-type mice, [29]. A total of $23,420$ genes (transcripts) are reported. There are $4,836$ transcripts which have zero count values for the $301$ cells. These transcripts were removed and the remaining dataset has $18,584$ rows.
Figure 1: Log scale eigenvalues of the trachea dataset
@@
Figure 1 shows the behavior of the $\log$ scale singular values. For this example, the first $87$ singular values capture $90\%$ of information and the first $159$ singular values capture $95\%$ of information. One can observe a fast decay for the first $20$ singular values and a slow decay for the others.
@@
Five methods available in the literature were used to show the performance of the proposed method. The calculations were performed on the same laptop computer equipped with i5-1135G7 processor, 8 GB of RAM, under Microsoft Windows 11 family and R environment version 4.4.1, [35]. The psvd R package of the proposed method is available from the comprehensive R archive network, http://cran.r-project.org/web/packages/psvd/. The classical SVD method is available through the R build-in svd() function which computes all singular values. The Lanczos algorithm based method for computing all singular values is available in the svd R package, [21]. A C++ library for large scale eigenvalue problem is available in the Spectra tool, https://github.com/yixuan/spectra. The RSpectra R package allows to use this library for computing the first $r$ singular values, [34]. The irlba R package, [2], also allows to compute the first $r$ singular values. The probabilistic SVD method is implemented in the rsvd R package, [10]. When possible, the largest $20$, $50$, $100$ or $150$ singular values were computed. Each method was run $10$ times using default settings and the average results were reported. For the proposed method, the following settings were used: $q=2$, tolerance ϵ=italic-$\epsilon=$1e-8 and $\eta=10$. The classical svd() function was used as reference and three parameters were examined: the runtime, percentage rate of reconstruction and the means squared error (MSE).
Table 3: Comparison results
Method runtime (s) reconst. Rate % MSE
$r=20$ svd 3.77 80.83 -
propack.svd 4.58 80.83 1.39e-8
RSpectra 0.53 80.83 1.53e-8
irlba 0.42 80.83 0.00160
rsvd 1.04 80.83 320.106
psvd 1.66 80.83 2.35e-8
$r=50$ svd 3.77 86.21 -
propack.svd 4.58 86.21 1.39e-8
RSpectra 1.29 86.21 1.56e-8
irlba 1.36 86.21 0.00554
rsvd 2.42 86.15 1740.22
psvd 2.61 86.21 7.77496
$r=100$ svd 3.77 91.11 -
propack.svd 4.58 91.11 1.39e-8
RSpectra 2.12 91.11 0.69e-8
irlba 5.42 91.11 0.00989
rsvd 4.92 91.01 2162.08
psvd 4.50 91.11 38.2471
$r=150$ svd 3.77 94.47 -
propack.svd 4.58 94.47 1.39e-8
RSpectra 3.18 94.47 1.51e-8
irlba 14.53 94.47 0.01732
rsvd 7.54 94.38 1688.06
psvd 6.93 94.47 46.3998
The results obtained are summarized in Table 3 where the proposed method appears in bold.
Runtime.
@@
All singular values are computed togheter using the classical svd and the propack.svd methods. The former method is slightly faster than the later. When the first $20$ are required, the RSpectra, irlba, rsvd and psvd methods use less runtime compare to the svd and propack.svd methods. For this example, when $r=150$, the RSpectra is the fastest followed by classical svd, propack.svd, psvd, rsvd and irlba. The runtime of this later method grows quickly when the number of the required singular values increases.
Reconstruction rate.
@@
Relation (3) was used to compute the reconstruction rate. The denominator is the same for all method and its value is obtained using the classical svd method results. The numerator is computed using the results of each method. The reconstruction rate is relatively the same for all methods. A very small difference appears for the rsvd method when $r\geq 50$.
Mean squared error.
@@
The differences between the singular values of each method and those obtained using the classical method where computed and used to calculate a mean squared error (MSE). The propack.svd and the RSpectra methods have quasi similar reconstruction rate as the classical svd method. The irlba method has near zero MSE value. The methods psvd and rsvd have more than zero MSE values. The psvd method MSE can be improved by using small tolerance $\epsilon$ value. The rsvd method has relatively higher MSE value and can lead to a significant gap between singular values, see Table 4 showing the last $10$ singular values for $r=100$. From the results of the rsvd method, we observed that the first $50$ estimates obtained using $r=100$ are more accurate than those obtained using $r=50$. The first $20$ singular values obtained by all methods are relatively accurate.
Table 4: $91$-th to $100$-th singular values
i svd psvd rsvd
91 10847.51 10846.65 10390.96
92 10804.19 10805.04 10331.54
93 10638.48 10638.33 10260.26
94 10600.40 10600.56 10040.75
95 10413.53 10413.28 9963.21
96 10376.98 10377.18 9920.82
97 10332.29 10332.33 9820.13
98 10289.21 10289.21 9691.56
99 10195.62 10195.62 9639.37
100 10002.21 10002.21 9432.25
The irlba and the rsvd methods cannot be used for small size matrices. Therefore, only RSpectra and the proposed psvd methods can be used to compute the $r$ largest singular values for any matrix.
3 Applications
In the above presentation, the eigendecomposition and the SVD consist of linear algebra manipulations of the matrix $\mathbf{X}$ entries. Truncating of the sum in (2) to $r<n$ leads to a $r$ rank approximation of the matrix $\mathbf{X}$. A direct application of the $r$ largest eigenvectors associated with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is the principal component analysis.

### v4

@@ 2.2 Relation with the power method
where $\lambda_{j}$ and $\gamma_{j}$ are the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$, respectively. Relation (23) shows that the smallest eigenvalue of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is one, meaning this matrix is always of full rank even for a rank deficient matrix $\mathbf{X}$. A safe computation of the norm $\|\mathbf{W}^{(t)}-\mathbf{W}^{(t-1)}\|_{\mathsf{F}}^{2}$, step 5 of alogithm 1, at any iteration requires to have the same number of columns for the matrix $\mathbf{W}$. When the matrix $\mathbf{X}$ is not of full rank, there exit a matrix $\mathbf{W}$ of size $n\times r$, $r<n$, allowing to have an exact reconstruction of the $\mathbf{X}$. Therefore, using alone the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ in algorithm 1 may lead to a problem for computing the above norm if we are not careful.
2.3 Convergence analysis and choice of the settings $\eta$ and $q$
@@
From the convex property of the objective function, the algorithm 1 is convergent. The rate of convergence depends on the behavior of the singular values of the matrix $\mathbf{X}$, the precision $\epsilon$ and the step parameter $\eta$. The rate of convergence for the probabilistic SVD is analyzed in [26, 25] where faster convergence is observed for matrix which singular values decay quickly. That means, larger the difference between successive singular values is, faster is the convergence. This result corresponds to the modulus conditions for the power method. Relation (12) shows a possibility to modify the singular values decay by using a power $q$ of the matrix in the EVD. From relation (23), a small $\eta$ value (near zero) may lead to near one eigenvalues for the matrix $\mathbf{G}$. On the other hand, a value greater to $1$ for the parameter $\eta$ leads to increasing the eigenvalues of the matrix $\mathbf{G}$. Therefore, the choices $q>1$ and $\eta\geq 1$, likely lead to rapid singular values decay for the matrix $\mathbf{G}$. However, when the matrix $\mathbf{X}$ has large eigenvalues, small $\eta$ value may be good.
@@
The matrix $\mathbf{G}=\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ which is computed once before starting the iterative search, may lead to numerical problem when $\mathbf{X}$ has some very small or high entries. At each iteration, a matrix multiplication and a Gram-Schmidt (GS) orthogonalization are performed. The modified GS orthogonalization method is summarized in algorithm 3, see Annexes. The matrices multiplication and the GS orthogonalization are the heavy computing time part of the proposed method which is implemented in the psvd R, [54], package (https://cran.r-project.org/web/packages/psvd/index.html). A C code is used to speed-up the matrices multiplication and the GS orthogonalization.
3 Experimental results
The implementation of algorithm 1 in the R package psvd has the following inputs: $X$ the matrix, $r$ (the first singular values, default is 2), $\eta$ (the gradient search positive scalar, default is 1), $q$ (power integer, default is 2), itmax (maximum iterations number, default is 200), errerr (convergence error default is 1e-8) and mySeed (a seed used for generating an initial solution, default is 50). The matrix $X$ is mandatory and $r$ is a user supply parameter. The choice of the parameters $\eta$ and $q$ are discussed in subsection 2.3. A large number of iterations number itmax may lead to accurate singular values, especially a large size data matrix where the convergence can be slow or for a challenging matrix example. The parameter err is used for convergence, see step 6 of algorithm 1. The parameter mySeed allows to have reproducible results for different runs on the same matrix.
To show the effectiveness of the proposed SVD method and to illustrate the choice of the settings, some numerical examples are used.
3.1 Small size matrix examples
@@
Three academic matrix examples, relations (24)-(25), and the Fisher’s iris data, [18], were used. The iris dataset corresponds to the length and the width measurements of the sepal and the petal for $3$ iris species: setosa, versicolor and virginica. $50$ measurements were observed for each species leading to $m=150$ and $n=4$.
$X_{a}=\left(\begin{array}[]{ccc}1&1&1\\ 0&2&1\\ 1&0&1\end{array}\right)\mbox{ ; }X_{b}=\left(\begin{array}[]{cccc}3&1&9&2\\ 10&4,&8&6\\ 7&6&12&1\\ 11&2&5&9\\ 1&1&1&0\end{array}\right)$ (24)
$X_{c}=\left(\begin{array}[]{ccccc}22&10&2&3&7\\ 14&7&10&0&8\\ -1&13&-1&-11&3\\ -3&-2&13&-2&4\\ 9&8&1&-2&4\\ 9&1&-7&5&-1\\ 2&-6&6&5&1\\ 4&5&0&-2&2\end{array}\right)$ (25)
@@
The singular values associates with these four examples are shown in Table 1. The second and the third eigenvalues for the matrix $\mathbf{X}_{c}$ are very close and the last two eigenvalues are zero. The matrix $\mathbf{X}_{c}$ is a challenging example for many SVD computation algorithms.
Table 1: Singular values for the matrices in relations (24) - (25) and for the iris dataset
@@
2 1.44504187 9.317337969 20.00000000 17.76103366
3 0.24697960 3.298813773 19.59591794 3.460930930
4 - 0.000000000 0.000000000 1.884826306
5 - - 0.000000000 -
Table 2: Iterations number for small size matrices examples, the maximum number is set to 200
@@
$\eta$ 0.01 0.1 1 2 5 10 100
$X_{a}$ 201 43 11 9 9 9 9
$X_{b}$ 24 9 6 6 6 6 6
$X_{c}$ 179 179 179 179 179 179 179
Iris 27 12 10 10 10 10 10
$q=3$
$\eta$ 0.01 0.1 1 2 5 10 100
$X_{a}$ 160 26 9 7 6 6 6
$X_{b}$ 8 5 5 5 5 5 5
$X_{c}$ 123 123 125 124 133 145 201
iris 10 8 8 7 7 8 7
$q=4$
$\eta$ 0.01 0.1 1 2 5 10 100
$X_{a}$ 85 16 7 6 5 7 7
$X_{b}$ 5 4 4 4 4 4 4
$X_{c}$ 140 178 201 201 201 201 201
iris 39 201 201 201 201 201 106
For a power parameter $q\in\{2,3,4\}$, the step parameter $\eta$ was varied. The number of iterations required for convergence, $8$ digits after the decimal point, are shown in Table 2 where $201$ means no convergence after $200$ iterations. For the examples $X_{a}$ and $X_{b}$, rapid convergence is observed by using $\eta\geq 1$ and $q\in\{2,3,4\}$. The settings $\eta\leq 0.1$ and $q=3$ seem appropriate for the example $X_{c}$ while $q=3$ and $\eta\geq 1$ are better for the iris dataset. Using $q=3$ or $q=4$ necessitates the same number of operations in the initialization step. Using $q=4$ may lead to huge entries in the matrix $\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$ for some data examples. Therefore, the choics $q=2,3$ are good compromise. The choice $q=3$ may be appropriate for large size dataset and it may be useful to scale the matrix entries before computing the SVD. The scaling of a nonnegative-valued matrix can consist of dividing each entry by the maximum value to have a matrix which entries vary in the interval $[0,1]$. The singular values are then divided by the same scalar.
3.2 High-dimensional example: trachea data
@@
The trachea dataset corresponds to gene expression profiling using the sequencing technology. The data was downloaded from the Gene Expression Omnibus, [5], using the access number GSE103354, file: GSE103354_Trachea_fullLength_TPM.txt.gz. This dataset is the full-length single cell RNA sequencing for $301$ trachea segments of C57BL/6 wild-type mice, [46]. A total of $23,420$ genes (transcripts) are reported. There are $4,836$ transcripts which have zero count values for the $301$ cells. These transcripts (rows) were removed and the remaining sparse dataset has $18,584$ rows, $58\%$ of zero-valued entries, few entries ($0.52\%$) of value greater than $1,000$ and a maximum entry value $609,926.3$.
Figure 1: Log scale singular values of the trachea dataset
@@
Figure 1 shows a $\log$ scale singular values behavior for the trachea dataset. There is a fast decay for the $20$ first singular values and a slow decay for the others. For this example, the $87$ and $159$ first largest singular values capture $90\%$ and $95\%$ of information, respectively.
@@
Five methods available in the literature were used to show the performance of the proposed method. The calculations were performed on the same laptop computer equipped with i5-1135G7 processor, 8 GB of RAM, under Microsoft Windows 11 family and R environment version 4.5.1, [54]. The psvd R package of the proposed method is available from the comprehensive R archive network, http://cran.r-project.org/web/packages/psvd/. The classical SVD method is available through the R build-in svd() function. The Lanczos algorithm based method for computing all singular values is available in the svd R package, [34]. A C++ library for large scale eigenvalue problem is available in the Spectra tool, https://github.com/yixuan/spectra. The RSpectra R package allows to use this library for computing the $r$ first singular values, [53]. The Lanczos bidiagonalization based irlba R package, [3], also allows to compute the $r$ first singular values. The probabilistic SVD method is implemented in the rsvd R package, [15]. The $20$, $50$, $100$ or $150$ first largest right singular vectors were computed for $5$ methods, see Table 3. The svd, propack.svd, RSpectra, and the psvd methods are almost deterministic and each leads to the same singular values for different runs on the same matrix. A small error, $\sim$1e-3, is observed for the irlba method singular values for different runs on the same matrix, this error is larger, ¿10, for the rsvd method. These later two methods were run $10$ times, the average results for the irlba method and the best run (minimum mean squares error) results for the psvd method are reported. Each method is run with the default settings. For the proposed method, the following settings were used: qq=2, tolerance $\epsilon$=1e-8, $\eta$=1 and itmax=200. The classical svd() function was used as reference and three parameters were examined: the runtime, percentage rate of reconstruction and the mean squared error (MSE). The MSE value is obtained by computing the differences between the $r$ first singular values of a given method and those of the classical SVD method. Then, the mean of the squares of the differences is computed and shown. The results obtained are summarized in Table 3 where the proposed method appears in bold.
Table 3: Comparison of the runtime, the reconstruction rate and the MSE for six SVD methods
$r$ Method runtime (s) reconst.rate % MSE
2020 svd 4.12 80.83 -
propack.svd 0.95 80.83 <<1.0e-10
RSpectra 0.51 80.83 <<1.0e-10
irlba 0.41 80.83 2.76e-8
rsvd 0.98 80.83 889.4
psvd 1.61 80.83 <<1.0e-10
$50$ svd 4.12 86.21 -
propack.svd 2.81 86.21 <<1.0e-10
RSpectra 1.27 86.21 <<1.0e-10
irlba 1.34 86.21 3.43e-7
rsvd 2.14 86.16 33,427.6
psvd 2.51 86.21 1.21
$100$ svd 4.12 91.11 -
propack.svd 3.75 91.11 <<1.0e-10
RSpectra 2.08 91.11 <<1.0e-10
irlba 5.42 91.11 4.03e-7
rsvd 4.51 91.01 36,076.5
psvd 4.37 91.11 14.63
$150$ svd 4.12 94.47 -
propack.svd 3.94 94.47 <<1.0e-10
RSpectra 3.12 94.47 <<1.0e-10
irlba 14.22 94.47 5.3e-3
rsvd 7.47 94.38 16,203.4
psvd 6.52 94.47 14.35
Runtime
@@
All singular values are simultaneously computed using the classical svd method. When the first $20$ are required, the RSpectra, irlba, rsvd, propack.svd and psvd methods use less runtime compare to the svd method. For this example, when $r=150$, the RSpectra is the fastest followed by classical svd, propack.svd, psvd, rsvd and irlba. The runtime of this later method grows quickly when the required number of the singular values increases.
Reconstruction rate
@@
Relation (3) was used to compute the reconstruction rate. The denominator is the same for all methods and its value is obtained using the classical svd method result. The numerator is computed using the result of each method. The reconstruction rate is almost the same for all methods. A very small difference appears for the rsvd method when $r\geq 50$.
Mean squared error
@@
The differences between the singular values of each method and those obtained using the classical method where computed and used to calculate the MSE. The propack.svd and the RSpectra methods have quasi similar reconstruction rate as the classical svd method. The first $20$ singular values obtained by all methods are relatively accurate. The irlba method has near zero MSE value. The psvd and rsvd methods have more than zero MSE values. The rsvd method has relatively higher MSE value and can lead to a significant gap between singular values, see Table 4 showing $10$ singular values when $r=100$ (a) or when $r=150$ (b).
Table 4: $91$-th to $100$-th singular values obtained using the classical svd, the proposed psvd and the rsvd methods: (a) $r=100$, (b) $r=150$
i svd psvd(a) rsvd(a) rsvd(b)
91 10,847.51 10,846.65 10,483.18 10,808.09
92 10,804.19 10,805.04 10,426.29 10,754.47
93 10,638.48 10,638.33 10,298.68 10,603.37
94 10,600.40 10,600.56 10,194.08 10,540.89
95 10,413.53 10,413.28 10,075.35 10,371.87
96 10,376.98 10,377.18 9,942.93 10,329.69
97 10,332.29 10,332.33 9,805.83 10,264.57
98 10,289.21 10,289.21 9,619.66 10,225.44
99 10,195.62 10,195.62 9,524.24 10,137.75
100 10,002.21 10,002.21 9,474.41 9,916.85
MSE - 0.161 231,365 3,206
Because of the random property of the rsvd method, the values shown in Table 4 for this method may be different for different runs while those for the classical method and the proposed methods remain unchanged.
The irlba and the rsvd methods should be used for only large size matrices. These $2$ methods and the RSpectra method are the fastest for computing a small number of first largest singular vectors. Only the propack.svd, the RSpectra and the proposed psvd methods can be used to compute the $r$ largest singular vectors for any matrix. For all methods in the comparison, only the $k$ first singular vectors of the proposed method are exactly a neural network autoencoder weight matrix columns. Despite its relatively long runtime, the proposed method may be useful to have a direct link between data singular vectors and the columns of a neural network weight matrix.
4 Application results
Truncating of the sum in relation (2) to $r<n$ leads to a $r$ rank matrix $\mathbf{X}$ approximation. The $r$ first largest eigenvectors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are used in the principal component analysis and the neural network autoencoder model.